# ENFORCE: Exact Nonlinear Constrained Learning with Adaptive-depth Neural Projection

### v2

@@ Abstract
Ensuring neural networks adhere to domain-specific constraints is crucial for addressing safety and ethical concerns while also enhancing prediction accuracy. Despite the nonlinear nature of most real-world tasks, existing methods are predominantly limited to affine or convex constraints. We introduce ENFORCE, a neural network architecture that guarantees predictions to satisfy nonlinear constraints exactly. ENFORCE is trained with standard unconstrained gradient-based optimizers (e.g., Adam) and leverages autodifferentiation and local neural projections to enforce any $\mathcal{C}^{1}$ constraint to arbitrary tolerance $\epsilon$. We build an adaptive-depth neural projection (AdaNP) module that dynamically adjusts its complexity to suit the specific problem and the required tolerance levels. ENFORCE guarantees satisfaction of equality constraints that are nonlinear in both inputs and outputs of the neural network with minimal (and adjustable) computational cost.
Keywords Constrained learning  $\cdot$ Hard-constrained neural networks  $\cdot$ Trustworthy AI  $\cdot$ Physics-informed machine learning
@@ 1 Introduction
Neural networks (NNs) are the backbone of many recent advancements in artificial intelligence (AI), excelling in tasks such as natural language processing, image analysis, and scientific discovery due to their modularity, simplicity, and strong generalization capabilities. However, their ability falls short when strict adherence to domain-specific constraints is required. Depending on the task, prior knowledge and rules about the system (e.g., from physics, safety, or ethics) are often available and are typically leveraged by humans in decision-making processes. In contrast, traditional data-driven methods such as NNs rely solely on data. Thus, trained NNs may be accurate on a training and test data set but still may not satisfy constraints exactly, leading to inconsistent predictions. This limitation not only generates substantial skepticism, hindering their adoption in real-world applications but can potentially result in erroneous or unfeasible outcomes when NNs are applied to decision-making tasks. Moreover, when prior domain knowledge is available and can be expressed as analytical expressions such as equations, ensuring that NNs adhere to this information is crucial to avoid the suboptimal utilization of valuable insights.
@@
Enforcing strict constraints in NNs is a promising area of research for many fields. For example, in AI for Science, integrating first-principle laws ensures physically consistent models enabling insightful scientific discovery [1, 2]. More broadly, constraining neural network predictions can have a transformative impact in domains where strict adherence to critical requirements is essential. This includes applications such as safety-critical systems (e.g., self-driving vehicles [3], healthcare [4]), fairness and bias mitigation [5, 6], and compliance with regulatory standards (e.g., in finance [7]). Additionally, with the rise of generative AI (GenAI), enforcing constraints on generation processes could mitigate potential risks and ensure trustworthy adherence to given criteria. Constraining the neural network output to adhere to strict rules is also beneficial to tackle traditional machine learning challenges, such as overfitting and underfitting in data-scarce regimes [8]. It is also plausible that such constraints could enable some level of extrapolation in the presence of incomplete data or in holes of convex hulls [9, 10]. Hence, the combination of enhanced accuracy and adherence to pre-existing domain knowledge could enable the safe use of NNs as surrogate models in computationally expensive applications [11].
@@
Enforcing constraints in NNs is not an easy task. Early approaches often relied on incorporating penalty terms into the loss function to reduce violations of physical laws [12]. Yet, these penalty-based methods offer no guarantees of constraint satisfaction (soft-constrained). In contrast, other methods aim to ensure strict adherence to analytical constraints by design (hard-constrained). For instance, one can use activation functions such as sigmoids or ReLU to bound outputs. To enforce analytical constraints beyond bounds, recent studies incorporate correction layers into NNs to project or complete predictions, ensuring they lie within the feasible region. Recent research has developed efficient methods to enforce analytical constraints defined by affine relationships between input and output variables or by convex regions (See Section 2). However, many real applications such as in science or sociology are inherently governed by nonlinear relationships between the involved variables [13, 14]. Existing approaches for handling nonlinear constraints predominantly rely on external root-finding algorithms or constrained optimization solvers (Section 2.2). These methods introduce significant computational overhead and complicate model construction, thereby compromising the modularity and flexibility typically associated with NNs.
@@
In this paper, we propose ENFORCE, a neural network architecture that ensures predictions strictly satisfy nonlinear constraints representing domain knowledge or critical requirements. ENFORCE is trained using standard unconstrained optimization techniques and leverages an adaptive-depth projection module to enforce constraints by construction without relying on external solvers.
@@ 2 Related work
This section reviews existing approaches to enforce constraints in NNs. According to the focus of this paper, a more in-depth analysis will be conducted on existing methods for enforcing algebraic equality constraints exactly.
@@ 2.1 Soft-constrained neural networks
One of the earliest approaches to embedding domain knowledge into NNs involves the use of soft constraints. Soft constraints are incorporated as penalty terms appended to the loss function, penalizing residuals of algebraic [15, 16] or differential equations underlying the system [17]. Physics-informed neural networks (PINNs) [12] represent a widely used framework designed to solve partial differential equations (PDEs) with NNs by employing soft constraints and collocation points. Although the soft-constrained approach places no restrictions on the complexity of the constraints, it has the drawback of not guaranteeing strict adherence to them. Furthermore, increasing the complexity of the loss function – especially when the different terms vary in nature or scale – can degrade the optimization performance of the neural network, often resulting in suboptimal accuracy [18, 19].
@@ 2.2 Hard-constrained neural networks
Hard-constrained NNs refer to methodological approaches ensuring that neural network predictions strictly adhere to analytical constraints. These constraints, explicitly encoded within the architecture, act as inductive biases, guiding the learning process toward compliance with domain knowledge or restrictions [20]. Architectures such as convolutional neural networks (CNNs) [21] and graph neural networks (GNNs) [22, 23] encode inductive biases by guaranteeing invariance with respect to patterns and symmetries. Simple analytical constraints can be enforced using differentiable functions, such as sigmoids or ReLU for output bounding and softmax for simplex constraints (non-negativity and normalization). Recent literature includes significant contributions for enforcing analytical inequality constraints, such as convex polytopes and convex sets more generally [24, 25, 26, 27, 28]. One can also constrain the neural network to guarantee specific functional characteristics, such as being 1-Lipschitz [29] or Lyapunov stable [30], nevertheless, this is not a core objective of this study. For a broad and recent review on hard-constrained NNs, the reader is also referred to [8].
Since this paper focuses on analytical equality constraints, the following literature review considers existing methods for this specific case.
@@
Many methods for encoding hard equality constraints utilize projection techniques, which correct preliminary neural network predictions by appending a non-trainable layer to the output. Projections can be formulated as optimization problems (e.g., distance minimization) or derived from geometric principles. For example, in [31] NN predictions of physical systems governed by PDEs are projected to ensure solutions satisfy the finite difference discretization of the underlying linear PDEs. A more general approach is the KKT-hPINN [32], which enforces linear equality constraints of the form $Ax+By=b$, though its applicability is limited to affine equations. Recently, HardNet [8] was introduced to enforce equality and inequality constraints affine in the output, without input restrictions, via a closed-form projection step. HardNet-Cvx [8] extends this approach to more general convex functions in the inputs, leveraging differentiable optimization layers supported by optimization solvers [33].
@@
Alternatively, the neural network can predict a partial set of output variables, $y_{P}\in\mathbb{R}^{N_{O}-N_{C}}$, and complete the prediction by solving the system of constraints based on this partial output. This approach ensures that the constraints are always satisfied. For instance, Beucler et al. introduced this concept to simulate physical systems such as climate modeling [34]. However, when the constraints are not explicitly defined, solving the system requires a root-finding solver. Similar approaches have been proposed within the hybrid modeling community, particularly in the serial configuration, where a fully data-driven method is used to predict certain inputs to a mechanistic model [35]. While studies like DC3 [25] have developed efficient backpropagation techniques, scenarios involving implicit nonlinear constraints can be computationally expensive to tackle with predict-and-complete methods. Moreover, the predict-and-complete approach can reduce regression accuracy by hindering the representation capabilities of NNs [34].
@@
One of the most popular fields of application of constrained learning is for solving (parametric) constrained optimization problems using NNs [36]. OptNet [37] is an optimization layer developed to solve quadratic programs. Angrawal et al. [33] expands the methodology to convex programs. They develop efficient differentiation techniques through such layers. However, the forward pass always requires the solution of a constrained optimization problem. Recently, Mukherjee and Bhattacharyya [38] brute-forced the constrained learning paradigm by training a neural network using an NLP solver such as IPOPT [39] instead of standard unconstrained optimization algorithms. Also, NNs have been formulated as constrained optimization problems and solved using deterministic global solvers [11]. However, these approaches pose severe limitations in terms of NNs and dataset size.
Other methods
Other methods have been proposed for constrained learning in NNs, mostly considering affine or convex regions. Many of them consider constraints only dependent on the input of the neural network [10, 27, 40, 41], others design strategies to include the dependence on both inputs and outputs [28, 42]. Recently, contributions to enforce general logic and linear constraints have been proposed by the neuro-symbolic AI community, developing constraining layers using logic programming to improve multi-label classification and synthetic data generation [43, 44].
However, to the best of our knowledge, there is no method designed to enforce any nonlinear equality constraints involving both the input and the output of the neural network by encoding them in the architecture and without relying on external solvers.
@@ 3 Preliminaries
Given a dataset $(x_{i},y_{i})_{i=1,...,N}$, we aim to train a neural network $f_{\theta}$ with parameters $\theta$ to approximate the underlying relationships while satisfying a set of known algebraic equality constraints $c(x,y)=0$. In general, $c$ can be any nonlinear function in the input $x$ and output $y$ of the neural network, incorporating domain knowledge or specifying critical requirements.
@@
i) The constraints $c(x,y)=0$ are feasible, ii) $N_{C}<N_{O}$, where $N_{C}$ is the number of equality constraints and $N_{O}$ is the output dimensionality of the neural network, i.e., there are available degrees of freedom to learn.
One way to enforce the neural network prediction $\hat{y}$ to satisfy the constraints is to project it onto the feasible hypersurface defined by $c(x,y)=0$. The projection can be defined as a mathematical program:
$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{T}W\|y-\hat{y}\|$ (1)
s.t. $\quad c(x,y)=0$
[…]
@@
When $c(x,y)$ is an affine function in the neural network input and output (i.e., $c(x,y)=Ax+By-b=0$), then the problem results in a quadratic program (QP) and a closed-form analytical solution is available for the KKT conditions. Based on such observation, Chen et al. proposed KKT-hPINN by stacking a fixed non-trainable projection layer on top of a neural network, to guarantee the prediction adheres to given affine functions [32]. However, the method is restricted to linear constraint as a closed-form is not available when $c$ is nonlinear.
We note that a closed-form is still available when generalizing to any function in the input $x$, as the projection operation is still a QP. We observe that, assuming an affine dependence on $y$ such that $c=C(x)y-v(x)-b=0$, $C(x)$ and $v(x)$ play the roles of linear coefficients and translation, respectively, in a local sense. That is, given a input $x_{i}$ and a prediction $\hat{y}_{i}$, then any function of $x_{i}$ is a constant term with respect to the optimization problem in Eq. 1, which becomes again a QP.
Enforcing affine functions is not new and is achieved also through other techniques [8, 40]. However, relaxing the assumption to allow for any nonlinear constraints of both the input and output of the neural network commonly results in decreased computational efficiency, as typically requires the use of constrained optimization techniques [38] or nonlinear solvers such as Newton’s methods [25].
@@ 4 Nonlinear constrained learning
We present ENFORCE, a framework designed for general and efficient nonlinear constrained learning. The method employs a computationally cheap adaptive neural projection and has no restriction on the nonlinearity of $\mathcal{C}^{1}$ constraints involving both the input and output vectors of the neural network. At inference time, the model predictions satisfy the constraints to arbitrary tolerance $\epsilon$.
@@ 4.1 AdaNP: Adaptive-depth neural projection
We reformulate the NLP in Eq. 1 and exploit the efficiency of quadratic projections to generalize the methodology to any nonlinear constraint. Assuming $c$ is of class $\mathcal{C}^{1}$, we use first-order Taylor expansion to locally linearize the constraints around the neural network input $x_{0}$ and prediction $\hat{y}$:
@@
where $J_{x}c$ and $J_{y}c$ are the Jacobian matrices with respect to the variable $x$ and $y$, respectively. Since the neural network input is fixed for a given sample, the approximation is exact in $x$, hence, $x=x_{0}$. Considering orthogonal projection for notation simplicity, the nonlinear optimization problem in Eq. 1 is locally approximated by a (linearly constrained) QP:
$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{2}$ (3)
s.t. $\quad c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$
Definition 1.
Given an input $x\in\mathbb{R}^{N_{I}}$ to a neural network $f_{\theta}$, its prediction $\hat{y}=f_{\theta}(x)\in\mathbb{R}^{N_{O}}$, and a set of constraints $c\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{N{{}_{C}}})$, with $N_{C}<N_{O}$, we define an operator $\mathcal{P}$ such that $\tilde{y}=\mathcal{P}(\hat{y})$ is the solution to the linearized quadratic program in Eq. 3, in the domain $\Omega$ where the constraints are defined.
In particular, $\tilde{y}=B^{*}\hat{y}+v^{*}$, with $B^{*}=I-B^{T}(BB^{T})^{-1}B$ and $v^{*}=B^{T}(BB^{T})^{-1}v$, where $I\in\mathbb{R}^{N_{O}\times N_{O}}$ is the identity matrix, and:
$B$ $=\left.J_{y}c\right|_{x,\hat{y}}$
$v$ $=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})$
Since a closed-form expression for the operator $\mathcal{P}$ is derived (Appendix A.1), we can define a differentiable neural projection (NP) layer representing the operator $\mathcal{P}$. The forward and backward passes of an NP layer are computationally cheap (more details on implementation and computational cost are given in Appendix A.2). However, the operator $\mathcal{P}$ projects the neural network prediction onto a linear approximation of the nonlinear constraints (i.e., the tangent hyperplane). The error that we introduce is proportional to the projection displacement $\ell_{p}=||\tilde{y}-\hat{y}||$. From this consideration, it follows that (1) the error is mitigated as the projection displacement is small, i.e., the neural network prediction is sufficiently accurate, and (2) a single NP layer may not ensure exact adherence to nonlinear constraints. Nevertheless, it is worth noting that a single NP layer guarantees strict satisfaction of equality constraints that are affine in $y$ and nonlinear in $x$, i.e., it efficiently enforces constraint classes considered in similar recent works [32, 8].
To address the challenge of ensuring exact adherence to nonlinear constraints, as highlighted in the second consideration above, we propose AdaNP: an adaptive-depth neural projection composition to enforce nonlinear constraint satisfaction to arbitrary tolerance $\epsilon$.
Definition 2.
@@
Then, given an arbitrarily small scalar $\epsilon$ and $n\in\mathbb{N}$ such that $|c(x,\tilde{y}_{n})|<\epsilon$, the feasible prediction $\tilde{y}_{n}$ is computed as:
$\tilde{y}_{n}=(\mathcal{P}_{1}\circ\dots\circ\mathcal{P}_{n})(\hat{y})$
AdaNP is a differentiable stack of $n$-NP layers that can be composed on every neural network backbone. The depth $n$ adjusts adaptively depending on the nonlinearities and the specified tolerance. Accurate NNs typically result in shallower AdaNP modules, since the projection displacement $\ell_{p}$ is related to the loss $\ell_{d}=||\hat{y}-y^{*}||$, where $y^{*}$ is the ground truth output. This introduces a trade-off between the complexity of the backbone and the required depth of AdaNP to satisfy the specified tolerance criteria. It is worth mentioning that a similar concept of iterative neural projections, although employed for a different application and objective, is also utilized in [45].
Analogy with Sequential Quadratic Programming
AdaNP can also be seen as an iterative method that recursively improves the solution of a linearized nonlinear program. Here, we notice the similarity to sequential quadratic programming (SQP) techniques. Specifically, AdaNP is a simple case of SQP method for which the objective function is naturally quadratic while the nonlinear constraints are linearized (in contrast to full SQP, in which the objective function is quadratically approximated). This observation could potentially allow to analyze the convergence of the method using SQP theory [46, 47].

### v3

# ENFORCE: Nonlinear Constrained Learning with Adaptive-depth Neural Projection

@@ Abstract
Ensuring neural networks adhere to domain-specific constraints is crucial for addressing safety and ethical concerns while also enhancing inference accuracy. Despite the nonlinear nature of most real-world tasks, existing methods are predominantly limited to affine or convex constraints. We introduce ENFORCE, a neural network architecture that uses an adaptive projection module (AdaNP) to enforce nonlinear equality constraints in the predictions. We prove that our projection mapping is 1-Lipschitz, making it well-suited for stable training. We evaluate ENFORCE on an illustrative regression task and for learning solutions to high-dimensional optimization problems in an unsupervised setting. The predictions of our new architecture satisfy $N_{C}$ equality constraints that are nonlinear in both the inputs and outputs of the neural network, while maintaining scalability with a tractable computational complexity of $\mathcal{O}(N_{C}^{3})$ at training and inference time.
@@ 1 Introduction
Neural networks (NNs) are the backbone of many recent advancements in artificial intelligence (AI), excelling in tasks such as natural language processing, image analysis, and scientific discovery due to their modularity, and strong generalization capabilities. However, their ability falls short when strict adherence to domain-specific constraints is required. Depending on the task, prior knowledge about the system (e.g., from physics, safety, or ethics) is often available and is typically leveraged by humans in decision-making processes. In contrast, data-driven methods such as NNs rely solely on data. Thus, trained NNs may be accurate on a training and test data set but still may not satisfy known constraints, leading to inconsistent predictions. This limitation not only generates substantial skepticism, hindering their adoption in real-world applications, but can also lead to erroneous or physically infeasible outcomes in decision-making processes. Moreover, when domain knowledge is available as analytical equations, ensuring that NNs adhere to this information is crucial to avoid a suboptimal utilization of expert insights and potentially reduce data demand [1].
@@
Enforcing strict constraints in NNs is a promising area of research for many fields. For example, in AI for Science, integrating first-principle laws ensures physically consistent models, enabling insightful scientific discovery [2, 3] or system modeling in engineering [4]. A prominent area of application involves using NNs to safely accelerate computationally intensive tasks such as learning surrogate models [5] or solutions to (parametric) constrained optimization problems, either in a supervised or unsupervised manner [6, 7, 8, 9]. More broadly, constraining neural network predictions can have a transformative impact in domains where strict adherence to critical requirements is essential, including safety-critical systems [10, 11], bias mitigation [12, 13], and compliance with regulatory standards [14]. Additionally, with the rise of generative AI (GenAI), enforcing constraints on generation processes could mitigate risks by ensuring that generated data respects given criteria [15]. Constraining the neural network output to adhere to strict rules is also beneficial to tackle traditional machine learning challenges, such as overfitting in data-scarce regimes [16].
@@
Enforcing constraints in NNs is not straightforward. The majority of existing approaches rely on incorporating penalty terms into the loss function to minimize constraint violations [17]. Yet, these penalty-based methods offer no guarantees of constraint satisfaction (soft-constrained). In contrast, other methods aim to ensure strict adherence to analytical constraints by design (hard-constrained). For instance, one can use sigmoid functions to bound outputs. To enforce analytical constraints, recent studies incorporate correction layers into NNs to project or complete the predictions, ensuring they lie within a feasible region. For example, developed methods can enforce constraints defined by affine relationships between input and output variables or by convex regions [18, 19, 16]. However, many applications, e.g., in science or sociology, are inherently governed by nonlinear constraints [20, 21]. Existing approaches for handling nonlinear constraints predominantly rely on external root-finding or constrained optimization solvers [7, 22]. These methods introduce significant computational overhead and complicate model development, thereby compromising the modularity and flexibility typically associated with NNs.
@@
In this paper, we propose ENFORCE, a neural network architecture that enforces predictions to satisfy nonlinear constraints. ENFORCE is trained using standard unconstrained optimization techniques and leverages an adaptive-depth neural projection (AdaNP) module to enforce constraints by construction without relying on external solvers.
@@ 2 Related work
This section reviews different approaches to enforcing constraints in NNs, with a focus on existing hard-constrained methods.
@@ 2.1 Soft-constrained neural networks
One of the earliest approaches to embedding domain knowledge into NNs involves the use of soft constraints. Soft constraints are incorporated as penalty terms appended to the loss function, penalizing residuals of algebraic [23, 24] or differential equations underlying the system [25]. Physics-informed Neural Networks (PINNs) [17] represent a widely used framework designed to solve partial differential equations (PDEs) with deep learning by employing soft constraints and collocation points. Although the soft-constrained approach places no restrictions on the complexity of the constraints, it has the drawback of not guaranteeing strict adherence to them. Furthermore, the complication of the loss landscape – especially when the different terms vary in nature or scale – can degrade the optimization performance of the neural network, often resulting in suboptimal accuracy [26, 27].
@@ 2.2 Hard-constrained neural networks
Hard-constrained neural networks refer to methodological approaches ensuring that neural network predictions adhere to given constraints by construction. Since this paper focuses on analytical equality constraints, the following literature review considers existing methods for this specific case.
@@
Many methods for encoding hard equality constraints utilize projection techniques, which correct preliminary neural network predictions by appending a non-trainable layer to the output. Projections can be formulated as optimization problems (i.e., distance minimization) or derived from geometric principles. For example, in [19] neural network predictions of physical systems governed by PDEs are projected to ensure solutions satisfy the finite difference discretization of the underlying linear PDEs. A more general approach is the KKT-hPINN [18], which enforces linear equality constraints in the inputs and outputs. Recently, HardNet [16] was introduced to enforce equality and inequality constraints affine in the output, without input restrictions, via a closed-form projection step.
@@
NNs can also predict a subset of output variables, $z\in\mathbb{R}^{N_{O}-N_{C}}$, and complete the prediction by solving the system of constraints based on this partial output (null-space methods). This approach ensures that the constraints are always satisfied. For instance, Beucler et al. introduced this concept to simulate physical systems such as climate modeling [28]. However, when the constraints are not available in explicit form, solving the system requires a root-finding solver. Similar approaches have been proposed within the hybrid modeling community, particularly in the serial configuration, where a fully data-driven method is used to predict unknown inputs to a mechanistic model [4]. While studies like DC3 [7] have developed efficient backpropagation techniques, scenarios involving implicit nonlinear constraints can be computationally expensive to tackle with predict-and-complete methods. Moreover, we rigorously show in Appendix B.6 that predict-and-complete approaches can suffer training instabilities if the constraints Jacobian is ill-conditioned [28].
@@
To enforce analytical constraints, researchers leveraged constrained optimization to deploy specialized layers or directly train the neural network. OptNet [29] is an optimization layer developed to solve quadratic programs. Angrawal et al. [30] expand the methodology to convex programs. They develop efficient differentiation techniques through such layers. Min et al. [16] leveraged such optimization layers to develop HardNet-Cvx, a neural network enforcing convex constraints. However, the forward pass always requires the solution of a constrained optimization problem. Recently, Mukherjee and Bhattacharyya [22] approached the constrained learning paradigm by training a neural network using a constrained optimization solver such as IPOPT [31] instead of standard unconstrained optimization algorithms. However, these approaches pose severe limitations in terms of neural network and dataset size.
Other methods have been proposed for constrained learning in NNs (cf. Appendix A). However, to the best of our knowledge, no existing method enforces nonlinear equality constraints involving both the input and output of a neural network by embedding them into the architecture while allowing training with unconstrained solvers.
@@ 3 Preliminaries
Given a dataset $(x^{*}_{i},y^{*}_{i})_{i=1,...,N}$, without loss of generality we consider a neural network $f_{\theta}$ with parameters $\theta$ to approximate the underlying relationships while satisfying a set of known algebraic equality constraints $c(x,y)=0$. In general, $c$ can be a nonlinear function in the input $x$ and output $y$ of the neural network, incorporating domain knowledge or specifying critical requirements.
@@
i) The constraints $c(x,y)=0$ are feasible and linearly independent, ii) $N_{C}<N_{O}$, where $N_{C}$ is the number of equality constraints and $N_{O}$ is the output dimensionality of the neural network, i.e., there are available degrees of freedom to learn.
One way to enforce the neural network prediction $\hat{y}$ to satisfy the constraints is to project it onto the feasible hypersurface (manifold) defined by $c(x,y)=0$. The projection operation can be defined as an optimization problem:
$\tilde{y}=\operatorname*{arg\,min}_{y}\frac{1}{2}(y-\hat{y})^{T}W(y-\hat{y})\quad\text{s.t.}\quad c(x,y)=0$ (1)
If $W$ is the identity matrix, the prediction is corrected by an orthogonal projection onto the feasible region. This can be interpreted as finding the feasible solution that minimizes the Euclidean distance from the original prediction $\hat{y}$. A local solution to the nonlinear program in Eq. 1 can be found by solving the first-order necessary optimality conditions, known as Karush–Kuhn–Tucker (KKT) conditions (cf. Chapter 12.3 in [33]). However, the latter is not necessarily straightforward as it may involve solving a system of nonlinear equations.
@@
When $c(x,y)$ is an affine function in the neural network input and output, then the problem results in a quadratic program (QP) and a closed-form analytical solution is available for the KKT conditions [18]. An extension to the closed-form is available when generalizing to any function in the input $x$, as the projection operation is still a QP. Consider an affine constraint on $y$ of the form $c=C(x)y-v(x)-b=0$, where $C(x)$ and $v(x)$ act as the linear coefficient matrix and translation vector, respectively. For a given input $x_{i}$ and prediction $\hat{y}_{i}$, any function of $x_{i}$ can be treated as constant with respect to the optimization problem in Eq. 1, which thus reduces to a QP.
Enforcing affine functions is not new and is also achieved through other techniques [16, 34]. However, relaxing the assumption to allow for nonlinear constraints of both the input and output of the neural network commonly results in decreased computational efficiency and stability, as it typically requires the use of constrained optimization [22] or root-finding solvers such as Newton’s methods [7].
@@ 4 Nonlinear constrained learning
We present ENFORCE, a framework designed for general and efficient nonlinear constrained learning. The method employs a computationally cheap adaptive neural projection module and has no restriction on the nonlinearity of $\mathcal{C}^{1}$ constraints involving both the input and output of the neural network. We prove the neural projection to be a $1$-Lipschitz mapping, implying adversarial robustness and stable gradient flow dynamics when compared to state-of-the-art constrained learning methods such as predict-and-complete.
@@ 4.1 AdaNP: Adaptive-depth neural projection
We locally approximate the nonlinear program in Eq. 1 and exploit the efficiency of quadratic projections to generalize the methodology to any nonlinear constraint. Assuming $c$ is of class $\mathcal{C}^{1}$, we use first-order Taylor expansion to locally linearize the constraints around the neural network input $x_{0}$ and prediction $\hat{y}$:
@@
where $J_{x}c$ and $J_{y}c$ are the Jacobian matrices with respect to the variable $x$ and $y$, respectively. Since the neural network input is fixed for a given sample, the linearization is exact in $x$, thus, $x=x_{0}$. Considering orthogonal projection, the nonlinear optimization problem in Eq. 1 is locally approximated by a (linearly constrained) QP:
$\tilde{y}=\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{2}\quad\text{s.t.}\quad c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$ (3)
Definition 1 (Projection operator $\mathcal{P}$).
Given an input $x\in\mathbb{R}^{N_{I}}$ to a neural network $f_{\theta}$, its prediction $\hat{y}=f_{\theta}(x)\in\mathbb{R}^{N_{O}}$, and a set of constraints $c\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{N{{}_{C}}})$, with $N_{C}<N_{O}$, we define an operator $\mathcal{P}$ such that $\tilde{y}=\mathcal{P}(\hat{y})$ is the solution to the linearized quadratic program in Eq. 3, in the domain $\Omega$ where the constraints are defined. In particular, $\tilde{y}=B^{*}\hat{y}+v^{*}$, with $B^{*}=I-B^{T}(BB^{T})^{-1}B$ and $v^{*}=B^{T}(BB^{T})^{-1}v$, where $I\in\mathbb{R}^{N_{O}\times N_{O}}$ is the identity matrix, $B=\left.J_{y}c\right|_{x,\hat{y}}$, and $v=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})$.
Given the closed-form expression of the operator $\mathcal{P}$ derived in Appendix B.1, we can define a differentiable neural projection (NP) layer representing the operator $\mathcal{P}$. The forward and backward passes of an NP layer are computationally cheap (more details on implementation and computational cost are given in Appendix D). However, the operator $\mathcal{P}$ projects the neural network prediction onto a linear approximation of the nonlinear constraints (i.e., the tangent hyperplane). The error that we introduce is proportional to the projection displacement $e_{D}=||\tilde{y}-\hat{y}||$. From this consideration, it follows that (1) the error is mitigated as the projection displacement is small, i.e., the neural network prediction is sufficiently accurate, and (2) a single NP layer cannot ensure exact adherence to nonlinear constraints. It is worth noting that a single NP layer guarantees strict satisfaction of equality constraints that are affine in $y$ and nonlinear in $x$, i.e., it efficiently enforces constraint classes considered in similar recent works [18, 16].
To address the challenge of satisfying nonlinear constraints, we propose AdaNP: an adaptive-depth neural projection composition that, under certain conditions, enforces nonlinear constraint satisfaction to arbitrary tolerance $\epsilon$.
Definition 2 (AdaNP module).
@@
Proposition 1 (Convergence of AdaNP).
Given an arbitrarily small scalar $\epsilon$, $n\in\mathbb{N}$ and assuming $\hat{y}$ in the positive reach (cf. Def. 4, Appendix B.4) of the constraints manifold =$\mathcal{M}=\{x\in\mathbb{R}^{N_{I}},\>y\in\mathbb{R}^{N_{O}}:c(x,y)=0\}$, then $\tilde{y}_{n}$ is computed as:
$\tilde{y}_{n}=(\mathcal{P}_{1}\circ\dots\circ\mathcal{P}_{n})(\hat{y}),$
and converges to a feasible prediction such that $|c(x,\tilde{y}_{n})|<\epsilon$ with linear convergence rate under constraint smoothness conditions (cf. proof in Appendix B.3).
AdaNP is a differentiable stack of $n$-NP layers that can be composed on every neural network backbone. The depth $n$ adjusts adaptively during training and inference depending on the nonlinearities and the specified tolerance (cf. Algorithm 1 in Appendix D.1 for details about the adaptive behavior). Accurate NNs typically result in shallower AdaNP modules, since the linearization error $e_{D}$ is related to the distance $||\hat{y}-y^{*}||$ between the neural network prediction $\hat{y}$ and ground truth output $y^{*}$. This introduces a trade-off between the complexity of the backbone and the required depth of AdaNP to satisfy the specified tolerance criteria.